OpenStudy (anonymous):

need help :( Use the associative law for addition and the distributive law to show that \(z(z_1 + z_2 + z_3) = zz_1 + zz_2 + zz_3\)

OpenStudy (helder_edwin):

the \(z\)'s are complex numbers?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

\[\large z(z_1+z_2+z_3)=z[z_1+(z_2+z_3)]=zz_1+z(z_2+z_3) \] \[\large =zz_1+(zz_2+zz_3)=zz_1+zz_2+zz_3 \]

OpenStudy (anonymous):

thanks !